If sec theta + ten theta = x , then sec theta equal to

Grade:11

1 Answers

THOTA HARSHITHA CHOWDARY
6 years ago
as (sectheta -tantheta)(sectheta+tantheta)=1 using we get an equ that sectheta-tantheta=1/x then on solving sectheta=1+x^2/2x